# SimDC: A High-Fidelity Device Simulation Platform for Device-Cloud Collaborative Computing

### v1

@@ Abstract
The advent of edge intelligence and escalating concerns for data privacy protection have sparked a surge of interest in device-cloud collaborative computing. Large-scale device deployments to validate prototype solutions are often prohibitively expensive and practically challenging, resulting in a pronounced demand for simulation tools that can emulate real-world scenarios. However, existing simulators predominantly rely solely on high-performance servers to emulate edge computing devices, overlooking (1) the discrepancies between virtual computing units and actual heterogeneous computing devices and (2) the simulation of device behaviors in real-world environments. In this paper, we propose a high-fidelity device simulation platform, called SimDC, which uses a hybrid heterogeneous resource and integrates high-performance servers and physical mobile phones. Utilizing this platform, developers can simulate numerous devices for functional testing cost-effectively and capture precise operational responses from varied real devices. To simulate real behaviors of heterogeneous devices, we offer a configurable device behavior traffic controller that dispatches results on devices to the cloud using a user-defined operation strategy. Comprehensive experiments on the public dataset show the effectiveness of our simulation platform and its great potential for application.
@@ I Introduction
Cross-device collaborative computing, such as federated learning [1], has emerged as a promising solution to meet the increasing demand for low-latency, high-throughput, and privacy-preserving applications [2, 3, 4, 5]. However, due to the large-scale deployment, heterogeneity of edge devices, and the diversity of their operational behaviors, the development and evaluation of device-cloud collaborative computing algorithms and applications face significant challenges [6, 7, 8]. Therefore, it is crucial to design and implement a high-fidelity device simulation platform for device-cloud collaborative computing, which provides a controlled and reproducible environment for real-world deployment. With this simulator, developers can quickly understand the operational effectiveness of their collaborative computing algorithms and applications under various conditions at a relatively low cost.
@@ II-A Device-cloud Collaborative Computing
Federated Learning (FL), as a typical device-cloud collaborative computing paradigm, has garnered considerable interest since its emergence due to its ability to compute collaboratively while keeping each device’s raw data stored locally [2, 3, 4]. FL operates as a distributed optimization challenge, where numerous clients, such as smartphones and IoT devices, collectively refine a global model. Here, each client harbors private data on-device and contributes to model training by retaining data proximity. The primary objective of FL is to determine model parameters that minimize the overall loss function across all devices. The process unfolds as follows: (1) Client Set: Define $\mathcal{K}$ to be the set of clients engaging in federated learning, where each client $k\in\mathcal{K}$ maintains a local dataset $\mathcal{D}_{k}$. (2) Global Model Parameters: Denote by $\mathbf{w}$ the vector of parameters for the global model. (3) Local Loss Function: For each client $k$, ascertain the local loss function, $F_{k}(\mathbf{w};\mathcal{D}_{k})$, to assess the performance of the model using parameters $\mathbf{w}$ on the local dataset $\mathcal{D}_{k}$. (4) Optimization Objective: The federated learning endeavor seeks to minimize the global loss function:
@@
where $p_{k}$ signifies the weight assigned to client $k$, under the constraint $\sum{k\in\mathcal{K}}p_{k}=1$. Typically, $p_{k}$ is proportionate to the size of the dataset held by client $k$. Throughout the training regimen, local devices relay their individually trained models to the cloud infrastructure for integration, utilizing aggregation strategies like FedAvg [1]. The merged global model is then redistributed to the local devices to facilitate successive rounds of local refinement. This cyclical interaction between local computation and global aggregation persists until the model attains convergence at round $T$.
@@ II-B Device-cloud Collaborative Computing Simulators
To tackle these challenges, pioneering work such as FS-Real[12, 13] embarked on an assessment of current device-cloud collaborative federated learning algorithms across various heterogeneous hardware platforms. This research revealed significant disparities between the performance of heterogeneous devices and the often assumed homogeneity in real settings. As a result, the authors introduced a logic-based simulation technique for federated learning that accounts for the heterogeneity of edge devices. Furthermore, FLEdge[20] enhances existing FL benchmarks by systematically evaluating client competencies, with particular emphasis on client behaviors and the impact of computational and communication bottlenecks. The researchers performed empirical studies on specialized hardware to emulate network traits and client conduct. Although these initiatives mark a foray into exploring the previously noted limitations, FS-Real still approaches the heterogeneity of edge devices using Android native simulation tools (which operate on HPCs), and does not capture the real-time reactions of actual devices. Meanwhile, FLEdge focuses chiefly on benchmarking assessments.
@@ III-A Task Design Specifications
In the SimDC platform, a task serves as the core operational unit, adhering to strict design specifications to ensure its validity and accuracy. Each task is assigned a unique identifier (task_id), allowing for precise identification and tracking to enhance efficiency in task management and monitoring. Users can simulate multiple devices with varying performance levels within a single task, all of which must execute the same computational process (operator flow) uniformly.
@@
Additionally, a task can request heterogeneous hybrid resources for computation, enhancing the adaptability and complexity of task execution. Each task can also be configured with a ”scheduling priority” parameter that determines the order in which tasks are executed if resources are sufficient. A higher priority value implies that the task is more likely to be executed ahead of others, which is crucial for time-sensitive or critical tasks.
@@ III-B Platform Architecture
The SimDC platform consists of several key modules that work together to implement its core functionality, as shown in Fig. 1, including:
@@ IV-A Heterogeneous Resources
Device Simulation has developed the PhoneMgr module to manage tasks on physical phone devices. PhoneMgr first handles the downloading and distribution of data, then employs Android Debug Bridge (ADB) commands to directly control the execution process of phone devices in the physical devices cluster. The physical devices cluster includes both local mobile phones and remote phones provided by the Mobile Service Platform (MSP). These devices vary in model and configuration, allowing users to categorize them based on customizable configurations, such as performance levels (e.g., High, Low) or specific device models. Notably, the physical devices cluster can designate certain phones for performance measurement during task execution, enabling the collection of runtime parameters to facilitate the development and tuning of edge algorithms.
@@ IV-B Hybrid Allocation Optimization
The SimDC platform allows tasks to be configured with hybrid heterogeneous resources for simulating devices. A task in Task Runner scheduled by the Task Scheduler comprises user’s algorithmic simulation requirements. Based on the designated total number of simulation devices and computing resources, Task Runner determines the optimal allocation of device quantities across hybrid heterogeneous resources, aiming to minimize the execution time.
@@ IV-C Physical Devices Performance Measurement
The SimDC platform has implemented a device management module called PhoneMgr, designed for efficient phone management of a physical devices cluster. PhoneMgr performs various operations and interface management for physical devices, primarily relying on ADB commands, and supports user-defined ADB scripts for phones running third-party applications (APKs). ADB is a versatile command-line tool capable of communicating with Android devices, executing numerous device actions such as installing and debugging applications, and providing access to the Unix shell. This client-server architecture tool effectively enhances the flexibility and control of device management.
@@ V-A DeviceFlow Architecture
In real-world practical applications, as illustrated in Fig. 3, physical phones often exhibit diverse behaviors due to factors such as timezones, environmental networks, user actions, and inherent variability. The behavior of these large-scale devices can significantly influence the computational performance of cloud services in device-cloud collaboration scenarios, potentially leading to issues such as model aggregation bias[15, 38]. To accurately simulate the operation of large-scale devices in such scenarios, we propose DeviceFlow, a device behavior traffic controller designed to emulate the interaction traffic between devices and cloud services.
When edge devices collaborate with cloud services, they typically upload computation results to storage spaces upon task completion and transmit messages to cloud services. Cloud services subsequently retrieve the corresponding data from storage based on the received messages for further processing. DeviceFlow operates as an intermediary component, bridging edge devices and cloud services by managing message transmission. From the perspective of edge devices, DeviceFlow functions as a proxy for the cloud, while from the viewpoint of cloud services, it serves as a representation of the edge devices.
@@ V-B Message Dispatching Strategy
In DeviceFlow, users can define various strategies to control message dispatching and simulate device dropouts, thereby enabling high-fidelity and flexible simulation of device behaviors. The dispatching strategies are categorized into two primary types: real-time accumulated dispatching and rule-based dispatching.
@@ VI Experiments
We conduct specific experiments to assess the efficacy of our proposed SimDC platform. The experimental settings and results are outlined as follows.
@@ VI-A1 Dataset and Model
Device-cloud collaboration deployment is one of the most common scenarios in recommendation systems. Within this study, we focus on one of the crucial problems in this area: click-through rate (CTR) prediction. We utilize the public Avazu dataset, which contains a vast amount of advertising click data, including features such as device information, application and website classification, and anonymized classification variables. For our experiments, we select a subset comprising approximately 2 million records, covering 100,000 unique devices identified by device_id as the training set, and we additionally select records from 1,000 unique devices as the test set.
@@ VI-A2 Experimental Environment
We have constructed two heterogeneous resource clusters for the simulation experiments: a Ray cluster and a physical devices cluster. The Ray cluster is used for large-scale logical simulations, with a default configuration of 200 CPU cores and 300 GB of memory, supporting elastic scaling. When utilizing the Ray cluster, each task creates a placement group, where each actor corresponds to a resource bundle, simulating different grades of devices. In this experiment, we categorize the simulated devices into two grades: High devices with a default configuration of 4 CPU cores and 12 GB of memory, and Low devices with 1 CPU core and 6 GB of memory. The resources configured to each actor can be adjusted based on actual needs to meet different experimental requirements.
In the physical devices cluster, we have a default configuration of 10 local physical devices and 20 remote MSP devices. Similar to the logical simulation, the physical devices are divided into High (4 devices, with more than 8 GB of memory) and Low (6 devices, with less than 8 GB of memory) grades. MSP devices are also categorized into High (13 devices) and Low (7 devices) grades. Physical devices can be further classified based on device models, CPU frequency, NPU support, etc., while the logical simulation can mirror such classifications by appropriately setting up resource bundles for the simulated devices.
Since current efforts to simulate large-scale device computations are primarily focused on federated learning frameworks, we were limited to comparing the effectiveness of the SimDC platform in large-scale device simulations with several existing mainstream federated learning frameworks, including FedScale[36] and FederatedScope[4]. It is important to note that SimDC fundamentally differs from these federated learning frameworks. SimDC is a device simulation platform designed for device-cloud collaborative scenarios, capable of simulating edge-side training, inference, specific calculations, and collective behavior traffic, while federated learning is just one of its applications. In related experiments, we focus solely on comparing the performance of these methods in large-scale device simulation calculations.
@@ VI-B1 Simultaneous Large-Scale Simulations and Physical Device Performance Measurement
Tasks can request and utilize hybrid heterogeneous resources for computation, enabling the collection of performance metrics from physical devices within the clusters while conducting large-scale logical simulations. In our experiment, we simulate 500 High and 500 Low grade devices. Additionally, we configure 5 physical devices per grade exclusively for training and performance measurement. These devices are not reused as computation units in a single round for simulating device computations.

### v2

@@ Abstract
The advent of edge intelligence and escalating concerns for data privacy protection have sparked a surge of interest in device-cloud collaborative computing. Large-scale device deployments to validate prototype solutions are often prohibitively expensive and practically challenging, resulting in a pronounced demand for simulation tools that can emulate real-world scenarios. However, existing simulators predominantly rely solely on high-performance servers to emulate edge computing devices, overlooking (1) the discrepancies between virtual computing units and actual heterogeneous computing devices and (2) the simulation of device behaviors in real-world environments. In this paper, we propose a high-fidelity device simulation platform, called SimDC, which uses a hybrid heterogeneous resource and integrates high-performance servers and physical mobile phones. Utilizing this platform, developers can simulate numerous devices for functional testing cost-effectively and capture precise operational responses from varied real devices. To simulate real behaviors of heterogeneous devices, we offer a configurable device behavior traffic controller that dispatches results on devices to the cloud using a user-defined operation strategy. Comprehensive experiments on the public dataset show the effectiveness of our simulation platform and its great potential for application. The code is available at https://github.com/opas-lab/olearning-sim.
@@ I Introduction
Cross-device collaborative computing, such as federated learning (FL) [1], has emerged as a promising solution to meet the increasing demand for low-latency, high-throughput, and privacy-preserving applications [2, 3, 4, 5]. However, due to the large-scale deployment, heterogeneity of edge devices, and the diversity of their operational behaviors, the development and evaluation of device-cloud collaborative computing algorithms and applications face significant challenges [6, 7, 8]. Therefore, it is crucial to design and implement a high-fidelity device simulation platform for device-cloud collaborative computing, which provides a controlled and reproducible environment for real-world deployment. With this simulator, developers can quickly understand the operational effectiveness of their collaborative computing algorithms and applications under various conditions at a relatively low cost.
@@ II-A Device-cloud Collaborative Computing
Federated Learning (FL), as a typical device-cloud collaborative computing paradigm, has garnered considerable interest since its emergence due to its ability to compute collaboratively while keeping each device’s raw data stored locally [2, 3, 4]. FL operates as a distributed optimization challenge, where numerous clients, such as smartphones and IoT devices, collectively refine a global model. Here, each client harbors private data on-device and contributes to model training by retaining data proximity. The primary objective of FL is to determine model parameters that minimize the overall loss function across all devices. The process unfolds as follows: (1) Client Set: Define $\mathcal{K}$ to be the set of clients engaging in federated learning, where each client $k\in\mathcal{K}$ maintains a local dataset $\mathcal{D}_{k}$. (2) Global Model Parameters: Denote by $\mathbf{w}$ the parameters for the global model. (3) Local Loss Function: For each client $k$, ascertain the local loss function, $F_{k}(\mathbf{w};\mathcal{D}_{k})$, to assess the performance of the model using parameters $\mathbf{w}$ on the local dataset $\mathcal{D}_{k}$. (4) Optimization Objective: The federated learning endeavor seeks to minimize the global loss function:
@@
where $p_{k}$ signifies the weight assigned to client $k$, under the constraint $\sum_{k\in\mathcal{K}}p_{k}=1$. Typically, $p_{k}$ is proportionate to the size of the dataset held by client $k$. Throughout the training regimen, local devices relay their individually trained models to the cloud infrastructure for integration, utilizing aggregation strategies like FedAvg [1]. The merged global model is then redistributed to the local devices to facilitate successive rounds of local refinement. This cyclical interaction between local computation and global aggregation persists until the model attains convergence at round $T$.
@@ II-B Device-cloud Collaborative Computing Simulators
To tackle these challenges, some works like FS-Real[12, 13] embarked on an assessment of current device-cloud collaborative algorithms across various heterogeneous hardware platforms. This research revealed significant disparities between the performance of heterogeneous devices and the often assumed homogeneity in real settings. As a result, authors introduced a logic-based simulation technique for FL that accounts for the heterogeneity of edge devices. Further, FLEdge[20] enhances existing FL benchmarks by systematically evaluating client competencies, with particular emphasis on client behaviors and the impact of computational and communication bottlenecks. Though these initiatives mark a foray into exploring the previously noted limitations, FS-Real still approaches the heterogeneity of edge devices using Android native simulation tools (which operate on HPCs), and does not capture the real-time reactions of actual devices. Meanwhile, FLEdge focuses chiefly on benchmarking assessments.
@@ III-A Task Design Specifications
In SimDC platform, a task serves as the core operational unit, adhering to strict design specifications to ensure its validity and accuracy. Each task is assigned a unique identifier (task_id), allowing for precise identification and tracking to enhance efficiency in task management and monitoring. Users can simulate multiple devices with varying performance levels within a single task, all of which must execute the same computational process (operator flow) uniformly.
@@
Additionally, a task can request heterogeneous hybrid resources for computation, enhancing the adaptability and complexity of task execution. Each task can also be configured with a ”scheduling priority” parameter that determines the order in which tasks are executed if resources are sufficient.
@@ III-B Platform Architecture
SimDC contains several key modules that work together to implement its core functionality, as shown in Fig. 1, including:
@@ IV-A Heterogeneous Resources
Device Simulation developed PhoneMgr module to manage tasks on physical phone devices. PhoneMgr first handles the downloading and distribution of data, then employs Android Debug Bridge (ADB) commands to directly control the execution process of phone devices in the physical devices cluster. The physical devices cluster includes local mobile phones and remote phones provided by the Mobile Service Platform (MSP). These devices vary in model and configuration, allowing users to categorize them based on customizable settings, such as performance levels (e.g., High, Low) or specific device models. Notably, the physical devices cluster can designate certain phones for performance measurement during task execution, enabling the collection of runtime parameters to facilitate the development and tuning of edge algorithms.
@@ IV-B Hybrid Allocation Optimization
SimDC allows tasks to be configured with hybrid heterogeneous resources for simulating devices. A task in Task Runner scheduled by Task Scheduler comprises user’s simulation requirements. Based on the total number of simulation devices and computing resources, Task Runner determines the optimal allocation of device quantities across hybrid heterogeneous resources, aiming to minimize the execution time.
@@ IV-C Physical Devices Performance Measurement
SimDC implements a device management module called PhoneMgr, designed for efficient management of a physical devices cluster. PhoneMgr performs various operations and interface management for physical devices, primarily relying on ADB commands, and supports user-defined ADB scripts for phones running third-party applications (APKs). ADB is a versatile command-line tool capable of communicating with Android devices, executing numerous device actions such as installing and debugging applications, and providing access to the Unix shell. This client-server architecture tool effectively enhances the flexibility and control of device management.
@@ V-A DeviceFlow Architecture
In real-world practical applications, as illustrated in Fig. 3, physical phones often exhibit diverse behaviors due to factors such as timezones, environmental networks, user actions, and inherent variability. The behavior of these large-scale devices can significantly influence the computational performance of cloud services in device-cloud collaboration scenarios, potentially leading to issues such as cloud computing (e.g., model agg) bias[15, 38]. To accurately simulate the operation of large-scale devices in such scenarios, we propose DeviceFlow, a device behavior traffic controller designed to emulate the interaction traffic between devices and cloud services.
When edge devices collaborate with cloud services, they typically upload computation results to storage upon task completion and transmit messages to cloud services. Cloud services then retrieve the corresponding data from storage based on the received messages for further processing. DeviceFlow operates as an intermediary component, bridging edge devices and cloud services by managing message transmission. From the perspective of edge devices, DeviceFlow functions as a proxy for the cloud, while from the viewpoint of cloud services, it serves as a representation of the edge devices.
@@ V-B Message Dispatching Strategy
In DeviceFlow, users can define various strategies to control message dispatching and simulate device dropouts, thereby enabling high-fidelity and flexible simulation of device behaviors. The dispatching strategies are categorized into real-time accumulated dispatching and rule-based dispatching.
@@ VI Experiments
We conduct specific experiments to assess the efficacy of our proposed SimDC platform.
@@ VI-A1 Dataset and Model
Device-cloud collaboration deployment is one of the most common scenarios in recommendation systems. Within this study, we focus on one of the crucial problems in this area: click-through rate (CTR) prediction. We use the public Avazu dataset, which contains a vast amount of advertising click data, including features such as device information, application and website classification, and anonymized classification variables. For our experiments, we select a subset comprising approximately 2 million records, covering 100,000 unique devices identified by device_id as the training set, and we additionally select records from 1,000 unique devices as the test set.
@@ VI-A2 Experimental Environment
We construct two heterogeneous resource clusters for simulation experiments: a Ray cluster and a physical devices cluster. The Ray cluster is used for large-scale logical simulations, with a default configuration of 200 CPU cores and 300 GB memory, supporting elastic scaling. When utilizing the Ray cluster, each task creates a placement group, where each actor corresponds to a resource bundle, simulating different grades of devices. Further, we categorize the simulated devices into two grades: High devices with a default configuration of 4 CPU cores and 12 GB memory, and Low devices with 1 CPU core and 6 GB memory. The resources configured to each actor can be adjusted based on actual needs to meet different experimental requirements.
In the physical devices cluster, we have a default configuration of 10 local physical devices and 20 remote MSP devices. Similar to the logical simulation, the physical devices are divided into High (4 devices, with more than 8 GB memory) and Low (6 devices, with less than 8 GB memory) grades. MSP devices are also categorized into High (13 devices) and Low (7 devices) grades. Physical devices can be further classified based on device models, CPU frequency, NPU support, etc., while the logical simulation can mirror such classifications by setting up resource bundles for the simulated devices.
Since current efforts to simulate large-scale device computations are mainly focused on federated learning frameworks, we limit to comparing the effectiveness of SimDC in large-scale device simulations with several mainstream federated learning frameworks, including FedScale[36] and FederatedScope[4] (FS-Real[12, 13] is one version of FederatedScope). It is important to note that SimDC fundamentally differs from these federated learning frameworks. SimDC is a device simulation platform designed for device-cloud collaborative scenarios, capable of simulating edge-side training, inference, specific calculations, and collective behavior traffic, while federated learning is just one of its applications. In related experiments, we focus solely on comparing the performance of these methods in large-scale device simulation calculations.
@@ VI-B1 Simultaneous Large-Scale Simulations and Physical Device Performance Measurement
Tasks can request and utilize hybrid heterogeneous resources for computation, enabling the collection of performance metrics from physical devices within the clusters while conducting large-scale logical simulations. In our experiment, we simulate 500 High and 500 Low grade devices. Additionally, we configure 5 physical devices per grade as benchmarking devices exclusively for training and performance measurement. These devices are not reused as computation units in a single round for simulating device computations.
@@ VII Conclusion
Acknowledgments
We gratefully acknowledge the strong support and insightful guidance from our supervisor, Jinsong Liao. We thank the OPPO team members for their dedicated contributions: Software Engineers Benchi Xie, Yunku Chen, Hui Wang, Xiangmou Qu, Hongtao Pang, Jin Rui, Chenglong Zhao, Chunying Zhang, Fengli Fan; and Testing Engineers Cheng Jin, Yang Feng, Ruxin Wang, Jingjing Dai, and Ji Chen. We express gratitude to the anonymous reviewers of the ICDCS conference for their constructive comments on this work.